# On the future cover of a sofic shift

### v2

@@ Abstract.
The paper contains a new proof of the theorem by Krieger which establishes the canonicity of the future cover of a sofic shift. In addition the paper describes a method to produce new canonical covers from a given one, resulting in canonical covers related to, but generally different from the future cover.
@@ 1. Introduction
commutes. Thus the subshift of finite type $X_{\mathbb{K}(Y)}$ and the factor map $L_{\mathbb{K}(Y)}$ are both conjugacy invariants for $Y$. The same is of course also true for any other cover which is canonical in the same way. It must be emphasized that also components of a canonical cover that are not of maximal entropy as well as the wandering parts of the cover, which often arise even when $Y$ is mixing, also give rise to conjugacy invariants for $Y$. This observation is the main motivation for the work and the results that are presented below.
@@
The main new result of the paper is a theorem which out of a given canonical cover produces a new one. To appreciate this, recall that Krieger proved the uniqueness of the conjugacy $\psi_{\mathbb{K}}$ in the diagram above. Without this additional requirement it is obvious that a canonical cover is not unique up to isomorphism; for example the disjoint union of two copies of the future cover will again be canonical in the sense used so far, but not with unique lifts of conjugacies. In the following we shall therefore refer to a construction of covers as weakly canonical when it gives lifts of conjugacies, but possibly lifts that are not unique, while a construction which gives rise to unique lifts of conjugacies, as the future cover does, will be called a strongly canonical cover. The main new result of the paper, Theorem 4.6, describes a method to produce a strongly canonical cover out of a given weakly canonical one.
An important point concerning the construction is that it produces new strongly canonical covers of sofic shifts. Up to now only two strongly canonical covers for sofic shifts have been known; at least when we for simplicity ignore the doubling coming from the passage from right-resolving to left-resolving covers that can be achieved by ’time-reversal’. Besides the future cover it is the minimal right-resolving cover, and only for irreducible sofic shifts. It is therefore interesting to compare the strongly canonical covers that can be produced using Theorem 4.6 with the future cover and, in the irreducible case, the minimal right-resolving cover. For this we introduce a method to find the future cover of a sofic shift starting from an arbitrary given presentation. This is based on a variation of the so-called subset construction described in §3.3 of [LM]. More specifically, it is shown that a restriction to a subset of the vertexes in the subset construction, applied to an arbitrary presentation of the sofic shift, gives rise to a labeled graph whose merged graph is isomorphic to the future cover. Thus the future cover is the factor obtained from a subgraph of the subset construction by identifying vertexes with the same follower set. This fact is used to produce the examples occurring in the paper, and makes it easy to show by examples that the new strongly canonical covers that can be obtained from Theorem 4.6 in general are different both from the future cover, and in the irreducible case, also the minimal right-resolving cover.
@@ 2.1. Setup and notation
Let $G$ be a finite directed graph with edges (or arrows) $E_{G}$ and vertexes $V_{G}$. For $e\in E_{G}$ let $s_{G}(e)\in V_{G}$ be the start vertex and $t_{G}(e)$ the terminal vertex of $e$. We shall always assume that $G$ does not contain sinks or sources; i.e. that $s_{G}^{-1}(v)\neq\emptyset$ and $t_{G}^{-1}(v)\neq\emptyset$ for all $v\in V_{G}$. We extend the definition of $t_{G}$ to finite and left-infinite paths in $G$, and the definition of $s_{G}$ to finite and right-infinite paths in $G$, in the obvious way. We denote the edge shift of $G$ by $X_{G}$. Thus $X_{G}$ consists of the bi-infinite sequences
@@
The first lemma deals with factor codes, right-resolving graphs and right-closing factor codes. We refer to page 18, Definition 3.3.1, and Definition 8.1.8 in [LM], respectively.
@@
Recode, first using (the proof of) Theorem 3.2.1 of [LM] to get a labeled graph $(H,L_{H})$ and a conjugacy $\psi_{0}:X\to X_{H}$ such that
@@ 2.2. Regular vertexes and rays
Assume that $x\in X_{H}$ is regular. Then $s_{H}(x_{k})\in V_{H}$ is regular for all $k\in\mathbb{Z}$.
@@ 2.3. The future cover and Krieger’s theorem
Let $x=(x_{i})_{i=-\infty}^{\infty}\in X_{\mathbb{K}(Y)}$ and consider an $N\in\mathbb{N}$. By definition $s_{\mathbb{K}(Y)}(x_{-N})=F(y)$ for some $y\in Y$ and
@@
$L_{\mathbb{K}(Y)}(x_{-N})L_{\mathbb{K}(Y)}(x_{-N+1})\cdots L_{\mathbb{K}(Y)}(x_{N}))\in\mathbb{W}(Y)$
@@
$(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is regular.
Proof.
Let $F(y)\in V_{\mathbb{K}(Y)}$ and consider the element $\alpha_{Y}(y)\in X_{\mathbb{K}(Y)}$. Then $t_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{(-\infty,-1]})=F(y)$. Note that $L_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{(-\infty,-1]})=y_{(-\infty,-1]}$, which implies that
$\left\{y^{\prime}\in Y[0,\infty):\ L_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{(-\infty,-1]})y^{\prime}\in Y\right\}=F(y).$
Thanks to (8) this is all we need to show, cf. (3).
Lemma 2.14.
@@
$\cdots<n_{3}<n_{2}<n_{1}<n_{0}<0$
and $s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{n_{j}})=s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{n_{0}})$ for all $j$. Let $q_{j}:=L_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{[n_{j},n_{0}-1]})=y_{[n_{j},n_{0}-1]}$ and note that $p^{j}:=q_{j}^{\infty}\in Y$. We claim that $F(\sigma^{n_{0}}(y))\subseteq F(p^{j})$ for all $j\in\mathbb{N}$. To see this, let $z\in F(\sigma^{n_{0}}(y))$. Since $F(\sigma^{n_{0}}(y))=s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{n_{j}})=s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{n_{0}})=F(\sigma^{n_{0}}(y))$ and $L_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{[n_{j},n_{0}-1]})=q_{j}$, it follows that $F(\sigma^{n_{0}}(y))=\frac{F(\sigma^{n_{0}}(y))}{q_{j}}$, and hence also that
@@
which implies that $p^{j}_{(-\infty,-1]}z\in Y$, i.e. $z\in F(p^{j})$. We claim next that $F(\sigma^{n_{0}}(y))=F(p^{l})$ for some $l\in\mathbb{N}$. By passing to a subsequence we may assume for this that $F(p^{j})=F(p^{1})$ for all $j$. If $F(p^{1})\backslash F(\sigma^{n_{0}}(y))\neq\emptyset$, there is an element $z\in Y[0,\infty)$ such that $y_{[n_{j},n_{0}-1]}z\in Y[0,\infty)$ for all $j>1$, but $y_{(-\infty,n_{0}-1]}z\notin Y$, which is impossible. This proves the claim. It follows that $\alpha_{Y}(y)_{[n_{0},-1]}$ is a path, labeled by $y_{[n_{0},-1]}$, which starts at $F(\sigma^{n_{0}}(y))=F(p^{l})=s_{\mathbb{K}(Y)}(\alpha_{Y}(p^{l})_{0})$ and ends at $s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{0})=F(y)=s_{\mathbb{K}(Y)}(x_{N})$. Set $p:=\sigma^{-n_{0}}(p^{l})$ and
@@
Let $v\in V_{G}$. Since $v$ is regular there is an $x\in X_{G}$ such that $t_{G}(x_{-\infty,-1]})=v$ and $f_{G}(v)=F(L_{G}(x))$. Define $\theta(v)\in V_{\mathbb{K}(Y)}$ such that
@@
Let $e\in E_{G}$. Choose $x\in X_{G}$ such that $t_{G}(x_{-\infty,-1]})=s_{G}(e)$ and $f_{G}(s_{G}(e))=F(L_{G}(x))$. Set $x^{\prime}_{i}=x_{i+1},\ i\leq-2$, $x^{\prime}_{-1}=e$ and let $x^{\prime}_{i},i\geq 0$, be arbitrary subject to the condition that $x^{\prime}\in X_{G}$. Then $t_{G}(x^{\prime}_{(-\infty,-1]})=t_{G}(e)$ and by using that $(G,L_{G})$ is right-resolving it follows that
@@
$=\left\{y\in Y[0,\infty):\ L_{G}(x^{\prime}_{-\infty,-1]})y\in Y\right\}.$
@@
commutes. Note that $\psi\circ L_{\mathbb{K}(Y)}$ is right-closing because $L_{\mathbb{K}(Y)}$ is right-resolving and $\psi$ is a conjugacy. By Lemma 2.1 there is therefore a right-resolving labeled graph $(G,L_{G})$ and a conjugacy $\phi:X_{\mathbb{K}(Y)}\to X_{G}$ such that
@@
commutes. All together we get a sliding block code $\phi^{\prime}:X_{\mathbb{K}(Y)}\to X_{\mathbb{K}(Z)}$ such that
@@
Uniqueness: If $\phi_{i}:\ X_{\mathbb{K}(Y)}\to X_{\mathbb{K}(Z)}$, are conjugacies such that $L_{\mathbb{K}(Z)}\circ\phi_{i}=\psi\circ L_{\mathbb{K}(Y)}$, $i=1,2$, then
@@
Let $x\in\mathcal{R}(L_{\mathbb{K}(Y)})$. Since $L_{\mathbb{K}(Z)}\circ\phi_{1}(x)=\psi\circ L_{\mathbb{K}(Y)}=L_{\mathbb{K}(Z)}\circ\phi_{2}(x)$ and $L_{\mathbb{K}(Z)}$ is injective on $\mathcal{R}(L_{\mathbb{K}(Z)})$ by Lemma 2.8, it follows that $\phi_{1}(x)=\phi_{2}(x)$. Since $\mathcal{R}(L_{\mathbb{K}(Y)})$ is dense in $X_{\mathbb{K}(Y)}$ by Corollary 2.17, it follows that $\phi_{1}=\phi_{2}$.
@@ 2.5. The terminal component in the future cover of an irreducible sofic shift
The material in this section is probably folklore to experts, but I haven’t found it in the literature.
Let $H$ be a finite directed graph without sinks or sources. A vertex $v\in V_{H}$ is recurrent when there is a path $\gamma$ in $H$ such that $s_{H}(\gamma)=t_{H}(\gamma)=v$, and it is called transient otherwise. Let $V^{r}_{H}$ denote the set of recurrent vertexes and $V^{t}_{H}$ the set of transient vertexes. Two vertexes $v,w\in V_{H}$ communicate, and we write $v\rightsquigarrow w$, when there is a path $\gamma$ in $H$ with $s_{H}(\gamma)=v$ and $t_{H}(\gamma)=w$. When $v,w\in V_{H}^{r},\ v\rightsquigarrow w$ and $w\rightsquigarrow v$, we write $v\sim w$. This an equivalence relation $\sim$ on $V_{H}^{r}$ and the elements $C\in V_{H}^{r}/\sim$ define a strongly connected subgraph $H_{C}$ of $H$ with $V_{H_{C}}=C$ and $E_{H_{C}}=\left\{e\in E_{H}:\ s_{H}(e),\ t_{H}(e)\in C\right\}$. The elements of $V_{H}^{r}/\sim$ and the corresponding graphs $H_{C},\ C\in V_{H}^{r}$, will be referred to as the components in $H$ and the irreducible subshifts $X_{C}:=X_{H_{C}},\ C\in V_{H}^{r}/\sim$, as the components in $X_{H}$.
@@
Let $(H,L_{H})$ be a right-resolving, regular and follower-separated labeled graph and $C$ a terminal component in $H$. For each $v\in C$ there is a finite path $\gamma$ in $H$ such that $t_{H}(\gamma)=v$ and $L_{H}(\gamma)$ is synchronizing for $L_{H}(X_{H})$.
@@
Let $v\in C$. Since $(H,L_{H})$ is regular there is an element $z\in X_{H}$ such that $t_{H}(z_{(-\infty,-1]})=v$ and $f_{H}(v)=F(L_{H}(z))$. By Lemma 2.20 there is an $N\in\mathbb{N}$ such that $F(L_{H}(z_{[-N,-1]}))=F(L_{H}(z))$. By Proposition 3.3.16 in [LM] there is a word $u\in\mathbb{W}(L_{H}(X_{H}))$ such that $L_{H}(z_{[-N,-1]})u$ is synchronizing for $L_{H}(X_{H})$. Since $u$ is a prefix of an element from $F(L_{H}(z))=f_{H}(v)$ there is a path $\gamma_{1}$ in $H$ such that $s_{H}(\gamma_{1})=v$ and $L_{H}(\gamma_{1})=u$. Note that $t_{H}(\gamma_{1})\in C$ since $C$ is a terminal component. Since $H_{C}$ is strongly connected there is a path $\gamma_{2}$ in $H_{C}$ such that $s_{H}(\gamma_{2})=t_{H}(\gamma_{1})$ and $t_{H}(\gamma_{2})=v$. Set $\gamma:=z_{[-N,-1]}\gamma_{1}\gamma_{2}$ and note that $L_{H}(\gamma)$ is synchronizing for $L_{H}(X_{H})$ because $L_{H}(z_{[-N,-1]})u$ is a prefix of $L_{H}(\gamma)$.
@@
Let $Y$ be an irreducible sofic shift space. There is only one terminal component $C$ in $\mathbb{K}(Y)$ and $(\mathbb{K}(Y)_{C},L_{\mathbb{K}(Y)})$ is isomorphic to the minimal right-resolving presentation of $Y$.
@@
Let $G$ and $H$ be directed graphs with unique terminal components, $C$ and $C^{\prime}$, respectively. Assume that $\chi:X_{G}\to X_{H}$ is a conjugacy. It follows that $\chi(X_{C})=X_{C^{\prime}}$.
@@
(Corollary (2.16) in [Kr1]) Let $\psi:Y\to Z$ be a conjugacy of irreducible sofic subshifts. Let $(G,L_{G})$ and $(H,L_{H})$ be the minimal right-resolving presentations of $Y$ and $Z$, respectively. It follows that there is unique conjugacy $\phi:X_{G}\to X_{H}$ such that
(10) ${X_{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{G}}$$\scriptstyle{\phi}$${X_{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{H}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Z}$

### v3

@@ Abstract.
The paper contains a new proof of the theorem by Krieger which establishes the canonicity of the future cover of a sofic shift. In addition the paper describes a method to produce a new canonical cover from a given one, resulting in a canonical cover related to, but generally different from the future cover.
@@ 1. Introduction
commutes. Thus the subshift of finite type $X_{\mathbb{K}(Y)}$ and the factor map $L_{\mathbb{K}(Y)}$ are both conjugacy invariants for $Y$. The same is of course also true for any other cover which is canonical in the same way. It must be emphasized that also components of a canonical cover that are not of maximal entropy as well as the wandering parts of the cover, which often arise even when $Y$ is mixing, also give rise to conjugacy invariants for $Y$. While any cover of a sofic shift is a window to its internal structure, a canonical cover guarantees that what you see is genuinely intrinsic. In this way canonical covers offer one of the few tools we have for the study of sofic shifts; in particular the structure that set them apart from subshifts of finite type. This observation is the main motivation for the work and the results that are presented below.
@@
The main new result of the paper is a construction which produces a new canonical cover out of a given one. To appreciate this, recall that Krieger proved the uniqueness of the conjugacy $\psi_{\mathbb{K}}$ in the diagram above. Without this additional requirement it is obvious that a canonical cover is not unique up to isomorphism; for example the disjoint union of two copies of the future cover will again be canonical in the sense used so far, but not with unique lifts of conjugacies. In the following we shall therefore refer to a construction of covers as weakly canonical when it gives lifts of conjugacies, but possibly lifts that are not unique, while a construction which gives rise to unique lifts of conjugacies, as the future cover does, will be called a strongly canonical cover. The main new result of the paper, summarized by Theorem 4.6, describes a method to produce a strongly canonical cover out of a given weakly canonical one.
Up to now only two strongly canonical covers for sofic shifts have been known; at least when we for simplicity ignore the doubling coming from the passage from right-resolving to left-resolving covers that can be achieved by ’time-reversal’. Besides the future cover it is the minimal right-resolving cover, and only for irreducible sofic shifts. With the construction in the present paper we have now at least four different ways to construct a strongly canonical cover of an irreducible sofic shift and at least two for a general sofic shift. We finish the paper by describing the results of the constructions for the even shift, showing that all four covers are mutually different in this case.
@@ 2.1. Setup and notation
Let $G$ be a finite directed graph with edges (or arrows) $E_{G}$ and vertexes $V_{G}$. For $e\in E_{G}$ let $s_{G}(e)\in V_{G}$ be the start vertex and $t_{G}(e)$ the terminal vertex of $e$. We shall always assume that $G$ does not contain sinks or sources; i.e. for every vertex $v$ there are arrows $e_{1},e_{2}\in E_{G}$ such that $t_{G}(e_{1})=v=s_{G}(e_{2})$. We extend the definition of $t_{G}$ to finite and left-infinite paths in $G$, and the definition of $s_{G}$ to finite and right-infinite paths in $G$, in the obvious way. We denote the edge shift of $G$ by $X_{G}$. Thus $X_{G}$ consists of the bi-infinite sequences
@@
The first lemma deals with factor codes, right-resolving graphs and right-closing factor codes. We refer to Definition 3.3.1 and Definition 8.1.8 in [LM], respectively. The book [LM] is the bible of symbolic dynamics and we shall sometimes use notions and results from [LM] without explicit reference.
@@
Recode, first using the proof of Theorem 3.2.1 in [LM] to get a labeled graph $(H,L_{H})$ and a conjugacy $\psi_{0}:X\to X_{H}$ such that
@@ 2.2. Regular vertexes and rays
Assume that $(H,L_{H})$ is right-resolving and that $x\in X_{H}$ is regular. Then $s_{H}(x_{k})\in V_{H}$ is regular for all $k\in\mathbb{Z}$.
@@ 2.3. The future cover and Krieger’s theorem
Let $x=(x_{i})_{i\in\mathbb{Z}}\in X_{\mathbb{K}(Y)}$ and consider an $N\in\mathbb{N}$. By definition $s_{\mathbb{K}(Y)}(x_{-N})=F(y)$ for some $y\in Y$ and
@@
$L_{\mathbb{K}(Y)}(x_{-N})L_{\mathbb{K}(Y)}(x_{-N+1})\cdots L_{\mathbb{K}(Y)}(x_{N})\in\mathbb{W}(Y)$
@@
Lemma 2.14.
$(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is regular.
Proof.
Let $y\in F(y)$. To see that $F(y)$ is a regular vertex in $\mathbb{K}(Y)$, recall that $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is right-resolving. Since $F(y)=s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{0})$ the desired conclusion follows therefore by combining Lemma 2.13 with Lemma 2.6. ∎
@@
$\cdots<n_{3}<n_{2}<n_{1}<n_{0}<N$
and $s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{n_{j}})=s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{n_{0}})$ for all $j$. Let $q_{j}:=L_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{[n_{j},n_{0}-1]})=y_{[n_{j},n_{0}-1]}$ and note that $p^{j}:=q_{j}^{\infty}\in Y$. We claim that $F(\sigma^{n_{0}}(y))\subseteq F(p^{j})$ for all $j\in\mathbb{N}$. To see this, let $z\in F(\sigma^{n_{0}}(y))$. Since $F(\sigma^{n_{0}}(y))=s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{n_{j}})=s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{n_{0}})$ and $L_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{[n_{j},n_{0}-1]})=q_{j}$, it follows that $F(\sigma^{n_{0}}(y))=\frac{F(\sigma^{n_{0}}(y))}{q_{j}}$, and hence also that
@@
which implies that $p^{j}_{(-\infty,-1]}z\in Y$, i.e. $z\in F(p^{j})$. We claim next that $F(\sigma^{n_{0}}(y))=F(p^{l})$ for some $l\in\mathbb{N}$. By passing to a subsequence we may assume for this that $F(p^{j})=F(p^{1})$ for all $j$. If $F(p^{1})\backslash F(\sigma^{n_{0}}(y))\neq\emptyset$, there is an element $z\in Y[0,\infty)$ such that $y_{[n_{j},n_{0}-1]}z\in Y[0,\infty)$ for all $j>1$, but $y_{(-\infty,n_{0}-1]}z\notin Y$, which is impossible. This proves the claim. It follows that $\alpha_{Y}(y)_{[n_{0},-1]}$ is a path in $\mathbb{K}(Y)$, labeled by $y_{[n_{0},-1]}$, which starts at $F(\sigma^{n_{0}}(y))=F(p^{l})=s_{\mathbb{K}(Y)}(\alpha_{Y}(p^{l})_{0})$ and ends at $s_{\mathbb{K}(Y)}(\alpha_{Y}(y)_{0})=F(y)=s_{\mathbb{K}(Y)}(x_{N})$. Set $p:=\sigma^{-n_{0}}(p^{l})$ and
@@
Let $v\in V_{G}$. Since $v$ is regular there is an $x\in X_{G}$ such that $t_{G}(x_{(-\infty,-1]})=v$ and $f_{G}(v)=F(L_{G}(x))$. Define $\theta(v)\in V_{\mathbb{K}(Y)}$ such that
@@
Let $e\in E_{G}$. Choose $x\in X_{G}$ such that $t_{G}(x_{(-\infty,-1]})=s_{G}(e)$ and $f_{G}(s_{G}(e))=F(L_{G}(x))$. Set $x^{\prime}_{i}=x_{i+1},\ i\leq-2$, $x^{\prime}_{-1}=e$ and let $x^{\prime}_{i},i\geq 0$, be arbitrary subject to the condition that $x^{\prime}\in X_{G}$. Then $t_{G}(x^{\prime}_{(-\infty,-1]})=t_{G}(e)$ and by using that $(G,L_{G})$ is right-resolving it follows that
@@
$=\left\{y\in Y[0,\infty):\ L_{G}(x^{\prime}_{(-\infty,-1]})y\in Y\right\}.$
@@
commutes. Note that $\psi\circ L_{\mathbb{K}(Y)}$ is right-closing because $(\mathbb{K}(Y),L_{\mathbb{K}(Y)})$ is right-resolving and $\psi$ is a conjugacy. By Lemma 2.1 there is therefore a right-resolving labeled graph $(G,L_{G})$ and a conjugacy $\phi:X_{\mathbb{K}(Y)}\to X_{G}$ such that
@@
commutes. All together we get a sliding block code $\phi^{\prime}:=\theta_{0}\circ\phi:X_{\mathbb{K}(Y)}\to X_{\mathbb{K}(Z)}$ such that
@@
Uniqueness: If $\phi_{i}:\ X_{\mathbb{K}(Y)}\to X_{\mathbb{K}(Z)}$, $i=1,2$, are conjugacies such that $L_{\mathbb{K}(Z)}\circ\phi_{i}=\psi\circ L_{\mathbb{K}(Y)}$, then
@@
Let $x\in\mathcal{R}(L_{\mathbb{K}(Y)})$. Since $L_{\mathbb{K}(Z)}\circ\phi_{1}(x)=\psi\circ L_{\mathbb{K}(Y)}(x)=L_{\mathbb{K}(Z)}\circ\phi_{2}(x)$ and $L_{\mathbb{K}(Z)}$ is injective on $\mathcal{R}(L_{\mathbb{K}(Z)})$ by Lemma 2.8, it follows that $\phi_{1}(x)=\phi_{2}(x)$. Since $\mathcal{R}(L_{\mathbb{K}(Y)})$ is dense in $X_{\mathbb{K}(Y)}$ by Corollary 2.17, it follows that $\phi_{1}=\phi_{2}$.
@@ 2.5. The terminal component in the future cover of an irreducible sofic shift
The material in this section is probably folklore to experts, but much of it I haven’t found in the literature.
Let $H$ be a finite directed graph without sinks or sources. A vertex $v\in V_{H}$ is recurrent when there is a path $\gamma$ in $H$ such that $s_{H}(\gamma)=t_{H}(\gamma)=v$, and it is called transient otherwise. Let $V^{r}_{H}$ denote the set of recurrent vertexes and $V^{t}_{H}$ the set of transient vertexes. Two vertexes $v,w\in V_{H}$ communicate, and we write $v\rightsquigarrow w$, when there is a path $\gamma$ in $H$ with $s_{H}(\gamma)=v$ and $t_{H}(\gamma)=w$. When $v,w\in V_{H}^{r},\ v\rightsquigarrow w$ and $w\rightsquigarrow v$, we write $v\sim w$. This an equivalence relation $\sim$ on $V_{H}^{r}$ and an element $C\in V_{H}^{r}/\sim$ defines a strongly connected subgraph $H_{C}$ of $H$ with $V_{H_{C}}=C$ and $E_{H_{C}}=\left\{e\in E_{H}:\ s_{H}(e),\ t_{H}(e)\in C\right\}$. The elements of $V_{H}^{r}/\sim$ and the corresponding graphs $H_{C},\ C\in V_{H}^{r}$, will be referred to as the components in $H$ and the irreducible subshifts $X_{C}:=X_{H_{C}},\ C\in V_{H}^{r}/\sim$, as the components of $X_{H}$.
@@
Consider then a labeled graph $(H,L_{H})$. A subset $U\subseteq V_{H}$ of vertexes in $H$ is a hereditary subset when $e\in E_{H},\ s_{H}(e)\in U\ \Rightarrow\ t_{H}(e)\in U$. When this holds we let $H_{U}$ denote the graph with $V_{H_{U}}:=U$ and $E_{H_{U}}:=\left\{e\in E_{H}:\ s_{H}(e)\in U\right\}$. By restriction the labeling $L_{H}$ defines a labeling of $H_{U}$ and we denote the resulting labeled graph by $(H_{U},L_{H})$. We call this a heritary labeled subgraph of $(H,L_{H})$. Note that a terminal component $C\subseteq V_{H}$ of $H$ is a hereditary subset.
@@
Let $(H,L_{H})$ be a right-resolving, regular and follower-separated labeled graph and $C$ a terminal component in $H$. For each $v\in C$ there is a finite path $\gamma$ in $H$ such that $t_{H}(\gamma)=v$ and such that every path in $H$ with label $L_{H}(\gamma)$ terminates at $v$. In particular, $L_{H}(\gamma)$ is synchronizing for $L_{H}(X_{H})$.
@@
Let $v\in C$. Since $(H,L_{H})$ is regular there is a ray $z\in X_{H}$ such that $t_{H}(z_{(-\infty,-1]})=v$ and $f_{H}(v)=F(L_{H}(z))$. By Lemma 2.20 there is an $N\in\mathbb{N}$ such that $F(L_{H}(z_{[-N,-1]}))=F(L_{H}(z))$. By Proposition 3.3.16 in [LM] there is a word $u\in\mathbb{W}(L_{H}(X_{H}))$ such that the paths in $H$ with label $L_{H}(z_{[-N,-1]})u$ all terminate at the same vertex. Since $u$ is a prefix of an element from $F(L_{H}(z))=f_{H}(v)$ there is a path $\gamma_{1}$ in $H$ such that $s_{H}(\gamma_{1})=v$ and $L_{H}(\gamma_{1})=u$. Note that $t_{H}(\gamma_{1})\in C$ since $C$ is a terminal component. Since $H_{C}$ is strongly connected there is a path $\gamma_{2}$ in $H_{C}$ such that $s_{H}(\gamma_{2})=t_{H}(\gamma_{1})$ and $t_{H}(\gamma_{2})=v$. Set $\gamma:=z_{[-N,-1]}\gamma_{1}\gamma_{2}$ and note that every path in $H$ with label $L_{H}(\gamma)$ must terminate at $v$ since $L_{H}(z_{[-N,-1]})u$ is a prefix of $L_{H}(\gamma)$. ∎
@@
Let $Y$ be an irreducible sofic shift. There is only one terminal component $C$ in $\mathbb{K}(Y)$ and $(\mathbb{K}(Y)_{C},L_{\mathbb{K}(Y)})$ is isomorphic to the minimal right-resolving presentation of $Y$.
@@
In the following we let $(\mathbb{F}(Y),L_{\mathbb{F}(Y)})$ be the minimal right-resolving presentation of an irreducible sofic shift $Y$, also known as the Fischer cover of $Y$. The conclusion of Proposition 2.22 can then be written as
$(\mathbb{F}(Y),L_{\mathbb{F}(Y)})=(\mathbb{K}(Y)_{C},L_{\mathbb{K}(Y)}),$
when $C$ denotes the terminal component of $\mathbb{K}(Y)$.
@@
Let $G$ and $H$ be finite directed graphs with unique terminal components, $C$ and $C^{\prime}$, respectively. Assume that $\chi:X_{G}\to X_{H}$ is a conjugacy. It follows that $\chi(X_{C})=X_{C^{\prime}}$.
@@
(Corollary (2.16) in [Kr1]) Let $\psi:Y\to Z$ be a conjugacy of irreducible sofic subshifts. There is unique conjugacy $\psi_{\mathbb{F}}:X_{\mathbb{F}(Y)}\to X_{\mathbb{F}(Z)}$ such that
(10) ${X_{\mathbb{F}(Y)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{\mathbb{F}(Y)}}$$\scriptstyle{\psi_{\mathbb{F}}}$${X_{\mathbb{F}(Z)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{L_{\mathbb{F}(Z)}}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi}$${Z}$